OpenStudy (anonymous):

A number decreased by -3 is equal to -40. Write an equation.

OpenStudy (anonymous):

let that number be x so : x - 3 = -40

OpenStudy (anonymous):

Let the number be x so x-3=-40

OpenStudy (amorfide):

you want to denote "a number" with an unknown we will use the letter x x -3 is equal to -40 x-3=-40

OpenStudy (anonymous):

x-3=-40